OpenStudy (anonymous):

7^3/7^2 Equals What?

OpenStudy (cwrw238):

use a property of exponents: x^a / x^b = x^(a-b)

OpenStudy (cwrw238):

just plug in the values in your question into the above formula and work it out x=7,a=3 and b=2

OpenStudy (anonymous):

So 343 / 49

OpenStudy (anonymous):

Then 7^(3-2)?

OpenStudy (cwrw238):

that is one way of doing it but an easier is to use the law of exponents 7^3 / 7^2 = 7^(3-2) = 7^1 = 7

OpenStudy (cwrw238):

ok?

OpenStudy (anonymous):

That's good, but I'm confused. Where did you get the 1 from?

OpenStudy (cwrw238):

3-2 = 1

OpenStudy (cwrw238):

easy

OpenStudy (anonymous):

Oh wow I feel stupid, but math has always been my weak spot. Thanks anyway.

OpenStudy (cwrw238):

yw